If you are 600 m from a better when you see him hit a ball, how long will you wait to hear the sound of the bat hitting the ball (at 20
1
Expert's answer
2021-10-11T11:01:13-0400

"S = 600m"

"\text {Light speed } v_l =3*10^8\frac{m}{s}"

"\text {Sound speed at }20\degree C v_s=343\frac{m}{s}"

"t =\frac{S}{v_l}+\frac{S}{v_s};\frac{S}{v_l}\approx 0;"

"t \approx \frac{S}{v_s}=\frac{600}{343}\approx 1.75s"

"\text{Answer:}1.75s"